# Flow-Modulated Scoring for Semantic-Aware Knowledge Graph Completion

### v2

@@ 3.2.3. Module Integration and Overall Scoring Process
(17) $s_{(h,t)}=\sigma\left(\left[m_{j-1}^{h},m_{j-1}^{t}\right]\cdot W_{j-1}+b_{j-1}\right).$
@@
Secondly, we employ the Condition Flow-Matching Module. Given the contextual condition $\boldsymbol{z}=(m_{j-1}^{h},m_{j-1}^{t})=(\boldsymbol{h}^{*},\boldsymbol{t}^{*})$, where the second symbol representation used here for the CFM module is for the sake of convenience in representation, this module predicts a dynamic vector field $\boldsymbol{v}_{\theta}(t,\boldsymbol{x})$. This vector field characterizes the instantaneous velocity and direction of the transformation from an initial state $\boldsymbol{h}^{*}$ (i.e., $\boldsymbol{x}$ at $t=0$) to a target state $\boldsymbol{t}^{*}$ (i.e., $\boldsymbol{x}$ at $t=1$). We then use this vector field to dynamically modulate the basic static score of the entity pair as follows:
(18) $s_{(h,t)}=s_{(h,t)}\odot v_{\theta}(t,\boldsymbol{x}).$
@@
(19) $p(r|h,t)=\text{SoftMax}\left(s_{(h,t)}\right).$
The primary objective for relation prediction is to minimize the loss between the predicted probabilities and the ground-truth relations over the training triplets $\mathcal{D}$:
(20) $\mathcal{L}_{\textit{{\text{pred}}}}(\Phi_{S},\theta)=\sum_{(h,r,t)\in\mathcal{D}}J(p(r|h,t),r),$
where $J(\cdot)$ is the cross-entropy loss function, and $\Phi_{S}$ represents all learnable parameters of the Semantic Context Learning module and the static scoring Linear.
To train the entire model effectively, we adopt a joint training strategy. The overall loss function $\mathcal{L}$ combines the relation prediction loss with the Condition Flow-Matching loss. The Condition Flow-Matching loss ensures that the learned vector field $\boldsymbol{v}_{\theta}(t,\boldsymbol{x})$ accurately approximates the desired conditional dynamics. The final training objective is to minimize the weighted sum of these losses:
(21) $\mathcal{L}(\Phi_{S},\theta)=\mathcal{L}_{\textit{\text{pred}}}(\Phi_{S},\theta)+\lambda\mathcal{L}_{cfm}(\theta),$
@@
It is worth noticing that the initial context representation derived from $\boldsymbol{h}^{*}$ and $\boldsymbol{t}^{*}$ plays a dual role: it directly contributes to the base score for relation prediction, and it also serves as the condition $\boldsymbol{z}$ for the flow-matching module, thereby influencing the dynamic modulation factor.
@@
9:     $s_{(h,t)}=s_{(h,t)}\odot v_{\theta}(t,\boldsymbol{\mathit{x}})$;
10:     $\mathcal{L}_{cfm}(\theta)=\mathbb{E}_{t,q(\boldsymbol{\mathit{z}}),p_{t}(\boldsymbol{\mathit{x}}\mid\boldsymbol{\mathit{z}})}\left\|v_{\theta}(t,\boldsymbol{\mathit{x}})-u_{t}(\boldsymbol{\mathit{x}}|\boldsymbol{\mathit{z}})\right\|_{2}^{2}$;
11:     $\mathcal{L}_{pred}(\Phi_{S},\theta)=\sum_{(h,r,t)\in\mathcal{D}}J(p(r|h,t),r)$;
@@
4. Experiment
@@
  • RQ4: How does FMS modulate the static scores of triplets to capture the dynamic nature of relations?
4.1. Transductive Experiment Settings
Datasets.
We conduct experiments on six knowledge graph datasets: (i) FB15K, derived from Freebase, a large-scale KG of general human knowledge; (ii) FB15K-237, a subset of FB15K where inverse relations are removed; (iii) WN18, containing conceptual-semantic and lexical relations among English words from WordNet; (iv) WN18RR, a subset of WN18 where inverse relations are removed; (v) NELL995, extracted from the 995th iteration of the NELL system, containing general knowledge; (vi) DDB14, collected from Disease Database, a medical database containing terminologies, concepts (diseases, symptoms, drugs), and their relationships. The statistics above are summarized in Table 1.
Baselines.
Based on the concept of Knowledge Graph Completion, we categorize all baseline methods into four groups: (i) Emb-based: TransE (Bordes et al., 2013), ComplEx (Trouillon et al., 2016), DistMult (Yang et al., 2015), RotatE (Sun et al., 2019), SimplE (Kazemi and Poole, 2018), QuatE (Zhang et al., 2019); (ii) Rule-based: DRUM (Sadeghian et al., 2019), and PTransE (Peng et al., 2022); (iii) GNN-Based: R-GCN (Schlichtkrull et al., 2018), Pathcon (Wang et al., 2021), and LASS (Shen et al., 2022) (iv) LLM-Based: KG-BERT (Yao et al., 2019), KGE-BERT (Nadkarni et al., 2021), and GilBERT (Nassiri et al., 2022)
Evaluation Metrics.
For each test triplet $(h,r,t)$, we construct the queries: $(h,?,t)$  [^1] Some related work formulates this as predicting missing tail (or head) entities given a head (or tail) entity and a relation. These two problems (entity prediction and relation prediction) are reducible to each other. For instance, a relation prediction model $\Phi(\cdot|h,t)$ can be converted to a tail prediction model $\Gamma(\cdot|h,r)=\operatorname{SoftMax}_{t}(\Phi(r|h,t))$, and vice versa. Given this equivalence, this work focuses on relation prediction., where the answers are $r$. Analogous to the evaluation method employed in link prediction, the mean Reciprocal Rank (MRR) and Hits@N are selected as evaluation metrics, with definitions consistent with previous research. Higher MRR and H@N scores signify superior model efficacy.
4.2. Inductive Experiment Settings
Datasets.
Following  (Zhang and Yao, 2022), we utilize a total of twelve subsets, encompassing four distinct versions each, which are derived from established datasets: WN18RR, FB15k237, and NELL-995. Each subset is uniquely characterized by a disparate split between the training and test sets, ensuring a robust and comprehensive evaluation schema. For a detailed discourse on the specific splits and a nuanced presentation of the statistical attributes of each subset, one can refer to the comprehensive descriptions available in  (Zhang and Yao, 2022).
Baselines.
For inductive learning, traditional Embedding-based methods are not suitable for the task due to their inability to generalize to unseen entities, we choose the following two classes of classical and effective models: (i) Rule-based: RuleN (Meilicke et al., 2018), NeuralLP (Yang et al., 2017), DRUM (Sadeghian et al., 2019); (ii) GNN-based: GraIL (Teru et al., 2020), PathCon (Wang et al., 2021).
Evaluation Metrics.
Table 1. Statistics of all datasets. $\mathbb{E}[d]$ and $\mathrm{Var}[d]$ are mean and variance of the node degree distribution, respectively.
@@ 4.2. Inductive Experiment Settings
Table 2. Transductive performance comparison on all six datasets. The results here are expressed as percentages. The best results are shown in bold while the second-best results are shown in the underline.
@@
Table 3. Inductive results (Hits@10). The best results are shown in bold while the second-best results are shown underlined.
@@
Table 4. Results by relation category on FB15k-237.
@@
Table 5. Ablation study on key components of FMS.
@@
4.3. Overall Performance Comparison: RQ1
To answer RQ1, we conducted extensive experiments comparing FMS with the strongest baseline we know of, both in terms of transductive and inductive learning.
The transductive results are summarized in Table 2. FMS shows remarkable improvement across all metrics on six datasets. Specifically, FMS achieves a 1.2% relative, 1.7% relative, 0.4% relative, 2.5% relative, 10.6% relative, and 1.8% relative increase in MRR scores over the best baselines on FB15K, FB15k-237, WN18, WN18RR, NELL995, and DDB14, respectively. These consistent and significant gains across diverse datasets, particularly the substantial improvement on NELL995, underscore the robustness and superior capability of our proposed FMS model. Furthermore, it is important to note that, compared to all kinds of models, FMS achieves better results for MRR scores on all datasets, establishing its state-of-the-art performance.
To further analyze the performance of FMS, we decomposed FB15k-237 based on relation categories similar to those defined in (Wang et al., 2014). We redefined the 1-N task by specifically isolating a test subset where each entity pair has two or more relation types. As shown in Table 4, FMS achieved overwhelmingly superior performance, indicating its successful capture of the dynamic nature of relations.
(a) Results with different Top-K.
(b) Results with different context aggregators.
(c) Results with different hops.
Figure 4. Hyper-parameters analysis of FMS on FB15k-237.
The inductive results are summarized in Table 3. FMS shows remarkable improvement across Hits@10 metrics on twelve datasets. It is noteworthy that while the optimal performance of other models typically ranges between 70% and 90% except PathCon, the FMS model almost consistently achieves between 98% and 99.9%, demonstrating its overwhelming advantage in inductive learning capabilities. Furthermore, FMS maintains near-perfect performance across different versions and splits of various datasets, indicating that the model possesses excellent stability and robustness to data variations.
Table 6. Number of parameters of all models on DDB14.
Model TransE DisMult RotatE QuatE PathCon FMS
Param 3.7M 3.7M 7.4M 14.7M 0.06M 0.02M
Table 7. Number of parameters of all models on FB15k-237.
Model TransE DisMult RotatE QuatE PathCon FMS
Param 5.9M 5.9M 11.7M 23.6M 1.67M 0.35M
4.4. Ablation Study: RQ2
4.4.1. Key Modules Ablation Study
@@
The ablation study results are presented in Table 5, leading to the following key conclusions: The Top-K selection mechanism exerts the most significant impact on the overall FMS performance, with its absence leading to a substantial decline in performance, particularly on the WN18RR dataset. This underscores Top-K’s pivotal role in filtering critical information and suppressing noise. Secondly, the energy-model inspired scoring function is also crucial for evaluating and ranking candidate answers, and its removal results in a noticeable loss in performance. While the Condition Flow-Matching Module has a comparatively smaller impact, it still provides a stable performance gain for the model. Therefore, the synergistic interaction of these components is a key factor enabling the FMS model to maintain its state-of-the-art performance.
4.4.2. Sensitivity to Key Hyperparameters
@@
Figure 4(a) shows how FMS performance varies with different Top-K values. The results demonstrate that a moderate K value (specifically 10 in this case) is critical. An overly small K provides insufficient context to leverage, while an excessively large K introduces increasing noise, potentially overwhelming the model with excessive contextual information and causing a drop in performance.
@@
Figure 4(b) then shows the performance of FMS’s Semantic Context Learning Module when using different context aggregators. It can be seen that our designed self-attention aggregator performs best, while the mean aggregator performs worst.
@@
Figure 4(c) illustrates the selection of hop count on the graph. The results indicate that FMS achieves competitive performance at the first hop and optimal performance at the second hop. Further increasing the hop count, however, leads to performance degradation. This phenomenon is likely attributable to the exponential increase in neighbor sampling associated with deeper hops, which consequently introduces more noise. Nevertheless, owing to the mechanisms of Top-k selection and attention aggregation, the model maintains its state-of-the-art performance.
4.5. Efficiency Analysis: RQ3

### v3

@@ 3.2.3 Overall Scoring Process and Joint Training
$s_{(h,t)}=\sigma\left(\left[m_{j-1}^{h},m_{j-1}^{t}\right]\cdot W_{j-1}+b_{j-1}\right).$ (19)
@@
Secondly, we employ the Condition Flow-Matching Module. Given the contextual condition $\bm{z}=(\bm{h}^{*},\bm{t}^{*})=(m_{j-1}^{h},m_{j-1}^{t})$, where the first symbol representation used here for the CFM module is for the sake of convenience in representation, this module predicts a dynamic vector field $\bm{v}_{\theta}(t,\bm{x})$. This vector field characterizes the instantaneous velocity and direction of the evolution from an initial state $\bm{h}^{*}$ (i.e., $\bm{x}$ at $t=0$) to a target state $\bm{t}^{*}$ (i.e., $\bm{x}$ at $t=1$). We then use this vector field to dynamically modulate the basic static score of the entity pair as follows:
$s_{(h,t)}=s_{(h,t)}\odot v_{\theta}(t,\bm{x}).$ (20)
@@
$p(r|h,t)=\text{SoftMax}\left(s_{(h,t)}\right).$ (21)
The primary objective for relation prediction is to minimize the cross-entropy loss between the predicted probabilities and the ground-truth relations over the training triplets $\mathcal{D}$. This is formulated as minimizing the negative log-likelihood of the true relations:
$\mathcal{L}_{\textit{{\text{pred}}}}(\Phi_{S},\theta)=-\sum_{(h,r,t)\in\mathcal{D}}\log(p(r|h,t)),$ (22)
where $\Phi_{S}$ represents all learnable parameters of the Semantic Context Learning module and the static scoring Linear.
To train the entire model effectively, we adopt a joint training strategy. The overall loss function $\mathcal{L}$ combines the relation prediction loss with the Condition Flow-Matching loss. The Condition Flow-Matching loss ensures that the learned vector field $\bm{v}_{\theta}(t,\bm{x})$ accurately approximates the desired conditional dynamics. The final training objective is to minimize the weighted sum of these losses:
$\mathcal{L}(\Phi_{S},\theta)=\mathcal{L}_{\textit{\text{pred}}}(\Phi_{S},\theta)+\lambda\mathcal{L}_{cfm}(\theta),$ (23)
@@
It is worth noticing that the initial context representation derived from $m^{h}_{j-1}(\bm{h}^{*})$ and $m^{t}_{j-1}(\bm{t}^{*})$ plays a dual role: it directly contributes to the base score for relation prediction, and it also serves as the condition $\bm{z}$ for the flow-matching module, thereby influencing the dynamic modulation factor.
@@
9:  $s_{(h,t)}=s_{(h,t)}\odot v_{\theta}(t,\bm{\mathit{x}})$;
10:  $\mathcal{L}_{cfm}(\theta)=\mathbb{E}_{t,q(\bm{\mathit{z}}),p_{t}(\bm{\mathit{x}}\mid\bm{\mathit{z}})}\left\|v_{\theta}(t,\bm{\mathit{x}})-u_{t}(\bm{\mathit{x}}|\bm{\mathit{z}})\right\|_{2}^{2}$;
11:  $\mathcal{L}_{pred}(\Phi_{S},\theta)=-\sum_{(h,r,t)\in\mathcal{D}}\log(p(r|h,t))$;
@@
To systematically validate the proposed Flow-Modulated Scoring (FMS) framework, our experimental evaluation is structured across two subsequent sections, each addressing a critical and complementary task within Knowledge Graph Completion (KGC).
Section 4 is dedicated to Relation Prediction, the primary focus of this work. This task is not only pivotal for enabling fine-grained reasoning in knowledge graphs but also serves as the principal validation for the core design philosophy of FMS—its ability to capture the contextual dependency and dynamic evolutionary nature of relational semantics. The objective of this chapter is therefore to demonstrate the superior performance and technical advancements of FMS through a rigorous and comprehensive comparison against a wide array of state-of-the-art models.
Building upon this foundation, Section 5 extends our assessment to the more fundamental and canonical task of Entity Prediction. While the FMS framework was conceived to address the profound complexities of relation modeling, its powerful representation learning and reasoning capabilities are hypothesized to possess broader applicability. Consequently, this chapter aims to verify the versatility and generalization power of the FMS framework. By adapting the model for entity prediction, we seek to establish that FMS is not merely a specialized solution for a single task but rather a versatile framework capable of empowering a spectrum of KGC challenges. This will comprehensively showcase its potential and practical utility as a novel paradigm for knowledge graph completion.
4 Relation Prediction Experiment
@@
  • RQ4: How does FMS modulate the static scores of triplets to capture rich contextual information shaping relation semantics?
  • RQ5: How does FMS modulate the static scores of triplets to capture the dynamic evolution of relations?
4.1 Transductive Experiment Settings
4.1.1 Datasets
We conduct experiments on six knowledge graph datasets: (i) FB15K [70], derived from Freebase, a large-scale KG of general human knowledge; (ii) FB15K-237 [70], a subset of FB15K where inverse relations are removed; (iii) WN18 [71], containing conceptual-semantic and lexical relations among English words from WordNet; (iv) WN18RR [71], a subset of WN18 where inverse relations are removed; (v) NELL995 [72], extracted from the 995th iteration of the NELL system, containing general knowledge; (vi) DDB14 [19], collected from Disease Database, a medical database containing terminologies, concepts (diseases, symptoms, drugs), and their relationships. The statistics above are summarized in Table I.
TABLE I: Statistics of all datasets. $\mathbb{E}[d]$ and $\mathrm{Var}[d]$ are mean and variance of the node degree distribution, respectively.
@@ 4.1.1 Datasets
4.1.2 Baselines
Based on the concept of Knowledge Graph Completion, we categorize all baseline methods into four groups: (i) Emb-based: TransE [20], ComplEx [40], DistMult [22], RotatE [14], SimplE [73], QuatE [74]; (ii) Rule-based: DRUM [75], and PTransE [76]; (iii) GNN-Based: R-GCN [77], Pathcon [19], LASS [78] and CBLiP [51]; (iv) LLM-Based: KG-BERT [29], KGE-BERT [79], and GilBERT [80]
4.1.3 Evaluation Metrics
To ensure a fair comparison, our training and evaluation setup follows that of PathCon [19]222 https://github.com/hyren/PathCon. For each test triplet $(h,r,t)$, we formulate a query (h,?,t)(h,?,t) [^3] Some related work formulates this as predicting missing tail (or head) entities given a head (or tail) entity and a relation. These two problems (entity prediction and relation prediction) are reducible to each other. For instance, a relation prediction model $\Phi(\cdot|h,t)$ can be converted to a tail prediction model $\Gamma(\cdot|h,r)=\operatorname{SoftMax}_{t}(\Phi(r|h,t))$, and vice versa. Given this equivalence, this work focuses on relation prediction. We also discuss the entity prediction task in section 5. where the model’s objective is to predict the correct relation $r$. The model is required to rank the ground-truth relation $r$ against all other relations in the knowledge graph, which serve as negative samples. For instance, a negative candidate would be $(h,r^{\prime},t)$, where $r^{\prime}\neq r$. Analogous to link prediction, we report Mean Reciprocal Rank (MRR) and Hits@N under the filtered setting, which removes any other known true relations between the entity pair $(h,t)$ from the candidate list before ranking. Higher scores for these metrics indicate better model performance.
TABLE II: Best hyper-parameter configurations for FMS. The order for dataset-specific values is: FB15K, FB15K-237, WN18, WN18RR, NL995, DDB14.
Hyper-parameter Value / Setting
$\lambda$ 1.2
Epoch 20
Dimension (dim) 64
L2 Regularization $1\times 10^{-7}$
Learning Rate $5\times 10^{-3}$
Neighbor Aggregation Attention
Temperature 0.95
Dataset-specific Parameters
Batch Size 128 / 128 / 128 / 128 / 128 / 64
Context Hops 2 / 2 / 3 / 3 / 2 / 2
Neighbor Samples 16 / 16 / 8 / 8 / 8 / 8
Num. Top-K 10 / 10 / 4 / 4 / 3 / 3
Num. Attention Heads 4 / 4 / 4 / 4 / 4 / 4
4.1.4 Implementation Details
All experiments are conducted on two NVIDIA 4090 GPUs. Our investigations revealed that FMS’s performance is highly sensitive to certain hyperparameters, particularly the number of context hops and the Top-K value, which depend on dataset characteristics. Consequently, to achieve optimal performance, we conducted a comprehensive hyperparameter search independently for each dataset across the following pre-defined spaces:
  • Dimension of hidden states: {16, 32, 64, 128};
  • Weight of L2 loss term: $10^{-8}$, $10^{-7}$, $10^{-6}$, $10^{-5}$};
  • Learning rate: {0.0005, 0.005, 0.01, 0.05, 0.1};
  • Number of context hops: {1, 2, 3, 4};
  • Number of neighbor samples: {8, 16, 24, 32};
  • Number of Top-K: {3, 4, 5, 8, 10}.
The best-performing configurations derived from this search are detailed in Table II. For result stability, each experiment was repeated three times, and we report the average performance. Our implementation is open-sourced and available on GitHub to support reproducibility. [^4] https://github.com/yuanwuyuan9/FMS
4.2 Inductive Experiment Settings
4.2.1 Datasets
Following  [26], we utilize a total of twelve subsets, encompassing four distinct versions each, which are derived from established datasets: WN18RR, FB15k237, and NELL-995. Each subset is uniquely characterized by a disparate split between the training and test sets, ensuring a robust and comprehensive evaluation schema. For a detailed discourse on the specific splits and a nuanced presentation of the statistical attributes of each subset, one can refer to the comprehensive descriptions available in  [26].
4.2.2 Baselines
For inductive learning, traditional Embedding-based methods are not suitable for the task due to their inability to generalize to unseen entities, we choose the following two classes of classical and effective models: (i) Rule-based: RuleN [81], NeuralLP [82], DRUM [75]; (ii) GNN-based: GraIL [17], PathCon [19].
4.2.3 Evaluation Metrics
TABLE III: Transductive performance comparison on all six datasets. The results here are expressed as percentages. The best results are shown in bold while the second-best results are shown in the underline.
@@ 4.2.3 Evaluation Metrics
TABLE IV: Inductive results (Hits@10). The best results are shown in bold while the second-best results are shown underlined.
@@
TABLE V: Results by relation category on FB15k-237.
@@
TABLE VI: Ablation study on key components of FMS.
@@
4.3 Overall Performance Comparison: RQ1
To answer RQ1, we conducted extensive experiments comparing FMS with the strongest baseline we know of, both in terms of transductive and inductive relation prediction.
The transductive results are summarized in Table III. FMS demonstrates remarkable and consistent improvements across all six datasets. Specifically, FMS achieves a 1.2% relative, 1.7% relative, 0.4% relative, 2.5% relative, 10.6% relative, and 1.8% relative increase in MRR scores over the best baselines on FB15K, FB15k-237, WN18, WN18RR, NELL995, and DDB14, respectively. These consistent and significant gains across diverse datasets, particularly the substantial improvement on NELL995, underscore the robustness and superior capability of our proposed FMS model. Furthermore, it is important to note that, compared to all kinds of models, FMS achieves better results for MRR scores on all datasets, establishing its state-of-the-art performance.
To further analyze the performance of FMS, we decomposed FB15k-237 based on relation categories similar to those defined in [21]. We redefined the 1-N task by specifically isolating a test subset where each entity pair has two or more relation types. As shown in Table V, FMS achieved overwhelmingly superior performance, indicating its successful capture of the dynamic nature of relations.
The inductive results are summarized in Table IV. FMS shows remarkable improvement across Hits@10 metrics on twelve datasets. It is noteworthy that while the optimal performance of other models typically ranges between 70% and 90% except PathCon, the FMS model almost consistently achieves between 98% and 99.9%, demonstrating its overwhelming advantage in inductive learning capabilities. Furthermore, FMS maintains near-perfect performance across different versions and splits of various datasets, indicating that the model possesses excellent stability and robustness to data variations.
4.4 Ablation Study: RQ2
4.4.1 Key Modules Ablation Study
@@
The ablation study results are presented in Table VI, leading to the following key conclusions: The Top-K selection mechanism exerts the most significant impact on the overall FMS performance, with its absence leading to a substantial decline in performance, particularly on the WN18RR dataset. This underscores Top-K’s pivotal role in filtering critical information and suppressing noise. Secondly, the energy-model inspired scoring function is also crucial for evaluating and ranking candidate answers, and its removal results in a noticeable loss in performance. While each component is crucial, their collective impact exceeds the sum of their parts. This highlights that the synergistic interaction of semantic context selection, energy-based scoring, and dynamic flow modulation is the key factor enabling the FMS model to maintain its state-of-the-art performance.
(a) Results with different Top-K.
(b) Results with different aggregators.
(c) Results with different hops.
Figure 4: Hyper-parameters analysis of FMS on FB15k-237.
4.4.2 Sensitivity to Key Hyperparameters
@@
Figure 4LABEL:sub@fig:sub_topk_analysis shows how FMS performance varies with different Top-K values. The results demonstrate that a moderate K value (specifically 10 in this case) is critical. An overly small K provides insufficient context to leverage, while an excessively large K introduces increasing noise, potentially overwhelming the model with excessive contextual information and causing a drop in performance.
@@
Figure 4LABEL:sub@fig:sub_aggregator_analysis then shows the performance of FMS’s Semantic Context Learning Module when using different context aggregators. It can be seen that our designed self-attention aggregator performs best, while the mean aggregator performs worst.
@@
Figure 4LABEL:sub@fig:sub_hops_analysis illustrates the selection of hop count on the graph. The results indicate that FMS achieves competitive performance at the first hop and optimal performance at the second hop. Further increasing the hop count, however, leads to performance degradation. This phenomenon is likely attributable to the exponential increase in neighbor sampling associated with deeper hops, which consequently introduces more noise. Nevertheless, owing to the mechanisms of Top-k selection and attention aggregation, the model maintains its state-of-the-art performance.
TABLE VII: Number of parameters of all models on DDB14.
Model TransE DisMult RotatE QuatE PathCon FMS
Param 3.7M 3.7M 7.4M 14.7M 0.06M 0.02M
TABLE VIII: Number of parameters of all models on FB15k-237.
Model TransE DisMult RotatE QuatE PathCon FMS
Param 5.9M 5.9M 11.7M 23.6M 1.67M 0.35M
(a) DDB14 (14 relation types)
(b) WN18RR (11 relation types)
Figure 5: The learned correlation between the contextual relations of head/tail entities and the predicted relations. Case studies are shown for the DDB14 dataset (a) and the WN18RR dataset (b).
4.5 Efficiency Analysis: RQ3